OpenStudy (anonymous):

Medal for help- Use the distributive Property to simplify the expression. (-1)(4-c) A. 4-c ****** My answer B.-4+c C. 4+c D.-4-c

OpenStudy (anonymous):

@shamil98

OpenStudy (zzr0ck3r):

\(a(b+c)= ab+ac\) so \((-1)(4-c)= (-1)4+(-1)(-c)=-4+c\)

OpenStudy (anonymous):

I need help check my questions

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

I will give a reward

OpenStudy (anonymous):

TY :)